# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

### v1

@@ 3.3.2. Refiner Agent $G_{\text{ref}}$
It operationalizes the analyzer’s insights into a concrete optimized prompt. Compared to a direct rewriting approach, which may introduce irrelevant or unjustified changes, the refiner explicitly conditions on the improvement report, producing refinements that are focused, justifiable, and consistent with diagnostic evidence. The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$.
3.4. Evaluation
[…]
(6) $\text{Score}(\hat{p},\hat{r})=\langle s_{help},s_{corr},s_{coh},s_{comp},s_{verb}\rangle,$
where each $s_{m}\in[0,4]$ for $m\in\mathcal{M}$. These scores are normalized into the $[0,1]$ range and then aggregated into a single composite value. [^2] This evaluation methodology is implemented in our setup in Section 4.2. Additionally, We complement automatic metrics with a human evaluation (Section 5.3) that tests (H1) the usefulness, accuracy, and consistency of the reasoning and (H2) whether optimized prompts preserve the original intent; 14 expert annotators, all data scientists or AI engineers, participated.
4. Experimental Design and Setup
4.1. Datasets
@@
We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2023, 2024a), which provide human-annotated prompt-response pairs for evaluating prompt optimization in Large Language Models (LLMs). Both datasets share the same annotation schema across five quality dimensions, but differ in size and design.
We specifically build the retrieval corpus from the HelpSteer2 training dataset, where each instance $\tau_{i}=(p_{i},r_{i},\mathcal{S}_{i})$ is augmented with the reasoning assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$ produced during our training pipeline. Evaluation is conducted on both HelpSteer1 and HelpSteer2 validation prompts.
HelpSteer1 contains 35.3k training and .79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0-4 scale. It served as the first large-scale benchmark for aligning model outputs with human feedback and provides broad coverage of prompt-response styles.
HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024a). We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field.
Table 1. Overview of HelpSteer1/2 Annotation Dimensions. Each scored on a 0–4 scale.
@@
Complexity Intellectual depth required to write the response (e.g., basic competency vs. domain expertise).
@@
Reasoning–asset generation (corpus construction).
Generation models for evaluation.
For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model LLaMA-3-8B-Instruct (Dubey et al., 2024) as the backbone models. These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$. Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
Retrieval.
We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text. For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^3] A detailed analysis of $k$ values and their effects on performance is in Section 5.1.1. and attach their paired responses and reasoning assets:
$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k}.$
4.2. Evaluation Method and Metrics
Judge model.
We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024b) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality. For any prompt-response pair, the judge model returns five HelpSteer-aligned scores (shown in Table 1):
Evaluation Metrics.
Each candidate pair $(p,r)$ is evaluated across the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$. We normalize scores by dividing by $4$, yielding $\tilde{s}_{m}=s_{m}/4\in[0,1]$. The overall quality score is then defined as the average over all metrics:
$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$
This produces a single composite value in the range $[0,1]$, reflecting the overall response quality.
4.3. Prompt Design
To guide each agent in MA-SAPO, we carefully designed specialized prompts tailored to their roles in both the training and test phases. [^4] For brevity, we show shortened prompts; full templates are in our code repository.
  1. (1)
    The training prompts (Metric Explainer, Diagnostician, and Action Synthesizer) focus on generating structured reasoning assets by interpreting evaluation scores, diagnosing weaknesses, and synthesizing actionable improvements.
  2. (2)
    The test prompts (Analyzer and Refiner) leverage these assets to contrast retrieved exemplars with the original prompt and iteratively refine it while preserving the user’s intent.
These prompts form the backbone of MA-SAPO’s reasoning-driven optimization pipeline, ensuring interpretability, controllability, and consistency across all stages. The prompt templates are provided below:
Prompt Templates for Different Agents in Training and Testing Phases Metric Explainer Agent Prompt You are an evaluation explainer. Given a user prompt, a model response, and five 0–4 scores (helpfulness, correctness, coherence, complexity, verbosity), write ONE cohesive paragraph (6–9 sentences) that explains WHY each score was assigned and HOW to improve the response. Rules: - Output must be a single paragraph in plain text (no lists, no JSON, no headings) (…) Diagnostician Agent Prompt You are a precise evaluation doctor for LLM outputs. Task: Write ONE cohesive paragraph that (1) diagnoses the main Strength and Weaknesses, (2) explains why they matter in terms of the five metrics, and (3) prescribes concrete fixes. Rules: - Evidence-first: rely only on the prompt + response + reasoning. - Mention each metric once (helpfulness, correctness, coherence, complexity, verbosity) with a brief, specific rationale (…) Output: One paragraph. Action Synthesizer Agent Prompt Mission: Based on the structured diagnosis (Diagnostician Agent), provide a single actionable prompt suggestion that guides improvements while preserving the original intent. Focus on highlighting strengths, addressing weaknesses, and suggesting … Task: Using the diagnosis, propose concise guidance that: 1. Maintains the core topic and original goal of the prompt. 2. Suggests ways to strengthen depth, clarity, and structure. 3. Recommends addition or adjustment (…) Analyzer Agent Prompt You are the Analyzer Agent. Task: (1) Independently evaluate the original prompt (strengths and weaknesses). (2) Analyze retrieved prompt–response–feature triples from three agents. (3) Integrate insights into a unified analysis. Rules: - Begin with direct evaluation of the original prompt. - Evaluate each agent’s output concisely (strengths/weaknesses) (…) Output: A structured but cohesive paragraph-length analysis. Refiner Agent Prompt You are the Refiner Agent. Task: Using Analyzer feedback, optimize the original prompt to resolve all issues while preserving intent. Rules: - Extract the core improvement points from the Analyzer. - Prioritize Action Synthesizer suggestions over others (…) Output: Final optimized prompt.
4.4. Baselines

### v2

@@ 3.3.2 Refiner agent $G_{\text{ref}}$
It operationalizes the analyzer’s insights into a concrete optimized prompt. Compared to a direct rewriting approach, which may introduce irrelevant or unjustified changes, the refiner explicitly conditions on the improvement report, producing refinements that are focused, justifiable, and consistent with diagnostic evidence. The final optimized prompt $\hat{p}$ is then used to generate a corresponding optimized response $\hat{r}$. [^1] The detailed prompts and specific role descriptions for agents are provided in Appendix H.
4 Experimental design and setup
4.1 Datasets
We conduct experiments on the HelpSteer datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs scored on five quality dimensions using a 0–4 scale. [^2] Detailed descriptions of the five annotation dimensions are provided in Table 4 in Appendix A. Both datasets follow the same annotation schema but differ in size and construction, as summarized in Table 1. We build the retrieval corpus solely from the HelpSteer2 training set, which offers higher-quality annotations, broader prompt diversity, and preference labels between responses that provide richer supervision for our Analyzer and Refiner agents. HelpSteer2 also serves as a standard benchmark for reward modeling. Evaluation is conducted on the validation splits of HelpSteer1 and HelpSteer2.
Dataset Train Val Multi-turn Preference Ann.
HelpSteer1 35.3k 1.79k × ×
HelpSteer2 20.3k 1.04k
Table 1: Statistics of the HelpSteer1/2 datasets. Both share the same annotation schema of five quality dimensions scored on a 0–4 scale.
4.2 Models used
Reasoning–asset generation (corpus construction).
Generation models for evaluation.
For downstream evaluation, we adopt a representative API model GPT-4o (Achiam et al., 2023), and an open-source model Llama3-8B-Instruct (Dubey et al., 2024) as the backbone models. These models are responsible for generating (i) optimized prompts $\hat{p}$ from the Refiner Agent and (ii) the corresponding optimized responses $\hat{r}$ conditioned on $\hat{p}$. Both models are configured with temperature =0=0 to ensure deterministic decoding, and all other hyperparameters remain at their defaults.
Judge model.
We employ the ArmoRM-Llama3-8B-v0.1 reward model (Wang et al., 2024a) as an interpretable multi-objective evaluator ($90\%$ benchmark accuracy), which automatically scores response quality. For any prompt-response pair, the judge model returns five HelpSteer-aligned scores: [^3] A detailed description of the ArmoRM model architecture, multi-objective training methodology, and alignment with HelpSteer dimensions is provided in Appendix C.
Retrieval.
We adopt a sparse retriever based on BM25 (Robertson et al., 2009) over prompt text. For each test prompt $p_{\text{test}}$, we retrieve the top-$k$ training prompts with $k=3$, which we identify as the optimal setting for retrieval [^4] A detailed analysis of $k$ values and their effects on performance is provided in Section 5.2.1. and attach their paired responses and reasoning assets:
$\{(p_{j},r_{j},\mathcal{R}_{j})\}_{j=1}^{k=3}.$ (6)
4.3 Evaluation metrics
We evaluate each candidate pair $(p,r)$ on the five HelpSteer dimensions $\mathcal{M}=\{\text{help},\text{corr},\text{coh},\text{comp},\text{verb}\}$, where each raw score $s_{m}\in[0,4]$. Scores are normalized as $\tilde{s}_{m}=s_{m}/4\in[0,1]$, and the overall quality score is computed as the average across all dimensions:
$\mathrm{Score}(p,r)=\tfrac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tfrac{s_{m}}{4}.$ (7)
This yields a single composite score in $[0,1]$ representing overall response quality. We also conduct a human evaluation with 17 expert annotators to assess (H1) the usefulness, accuracy, and consistency of the reasoning, and (H2) whether optimized prompts preserve the original intent. [^5] Human evaluation results are reported in Section 5.3, with detailed protocols and per-dimension scores in Appendix F.
4.4 Baselines
We compare our method with six baselines in three categories: single-pass prompt optimization without retrieval (Direct Generation, CoT (Wei et al., 2022), Role Assignment (Wu et al., 2023)), retrieval-augmented optimization (RAG (Lewis et al., 2020)), and multi-agent optimization (MAD (Du et al., 2024), MARS (Zhang et al., 2025)). For all methods, the model first generates an optimized prompt and then produces the final response using the same backbone LLM. We evaluate all baselines with GPT-4o and Llama3-8B-Instruct, using temperature 0 and default decoding settings. For RAG, we retrieve the top-$10$ training examples with BM25; for MAD, we use $2$ agents and $3$ debate rounds; and for MARS, we use 3344 planning steps and $5$ refinement iterations. [^6] Baseline details are provided in Appendix B.
5 Experimental results and analysis
5.1 Main results
Methods GPT-4o Llama3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1
Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
RAG (sparse, k=10k{=}10) (2020) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
Chain-of-Thought (CoT) (2022) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
Role Assignment (2023) 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
MAD (2024) 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
MARS (2025) 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
HelpSteer2
Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
RAG (sparse, k=10k{=}10) (2020) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
Chain-of-Thought (CoT) (2022) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Role Assignment (2023) 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
MAD (2024) 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
MARS (2025) 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
Table 2: Main results on HelpSteer1 and HelpSteer2. Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg). Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks. For each dataset and backbone block, the best value in each metric column is shown in bold, and the second-best is underlined.
In our main experiments, we compare MA-SAPO with three groups of baselines: (1) single-pass prompting methods (Direct Generation, CoT, and Role Assignment), (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks (MAD and MARS). As shown in Table 2, MA-SAPO consistently achieves the best results across both HelpSteer1 and HelpSteer2 under both GPT-4o and Llama3-8B. Single-pass methods remain limited because they revise prompts without diagnostic grounding, often leading to brittle or inconsistent improvements. Role Assignment provides more explicit guidance, but still depends heavily on predefined roles. RAG benefits from retrieved exemplars, but without explicit reasoning about their usefulness, its refinements remain heuristic and less reliable.
Figure 2: KDE score distributions of Direct Generation and MA-SAPO across five HelpSteer dimensions and their average, aggregated over GPT-4o and Llama3-8B. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution. MA-SAPO consistently shifts the score distribution toward higher values.
Methods GPT-4o Llama3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1
$k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
MA-SAPO (kk=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2
$k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
MA-SAPO (kk=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
Table 3: Ablation results over agent number $k$ and agent-combination testing on HelpSteer1 and HelpSteer2. Best scores are shown in bold and second-best scores are underlined.
Multi-agent baselines such as MAD and MARS achieve stronger performance by coordinating multiple roles, but they still rely on repeated interaction cycles and treat evaluation mainly as a scalar target. In contrast, MA-SAPO converts evaluation outcomes into reusable reasoning assets and uses them to support evidence-backed refinement through the Analyzer-to-Refiner pipeline, yielding stronger and more interpretable improvements. For example, on HelpSteer1 with GPT-4o, MA-SAPO achieves an average score of 0.6486, outperforming MARS (0.5096) and RAG (0.4784). Similar trends hold across HelpSteer2 and Llama3-8B, and Figure 2 further shows that these gains reflect a consistent shift toward higher scores across dimensions rather than a few outlier cases. [^7] Per-backbone KDE distributions are provided in Appendix E. This suggests that MA-SAPO improves multiple quality aspects simultaneously, rather than over-optimizing a single metric.
5.2 Ablation study
5.2.1 Top-k retrieval variations
To further analyze the robustness of MA-SAPO, we examine two key design choices: retrieval depth and test-phase agent composition. First, to study how the number of retrieved training examples affects performance, we vary the retrieval depth $k\in\{1,2,3,4\}$. Table 3 reports results for both GPT-4o and Llama3-8B across HelpSteer1 and HelpSteer2. Across both backbones, increasing $k$ from 1 to 3 generally improves performance, with $k=3$ achieving the best average balance across the five metrics. When $k=1$, the Analyzer is limited to a single exemplar, restricting the diversity of available reasoning assets. Increasing $k$ broadens the evidence base, but $k=4$ shows slight declines in some dimensions, suggesting that retrieving too many exemplars can introduce noise and conflicting signals. Overall, these results indicate a clear trade-off between contextual diversity and diagnostic noise, with $k=3$ providing the most reliable setting in our experiments.
5.2.2 Test agents combination
We subsequently assess the effect of test-phase agent modularity by merging the two agents into a single combined agent, denoted as Test Agents Combination. As shown in Table 3, this variant remains competitive for both GPT-4o and Llama3-8B on HelpSteer1 and HelpSteer2, and occasionally surpasses the full framework on individual metrics. However, its average performance consistently falls below that of the full MA-SAPO, indicating that it cannot reliably reproduce the benefits of the structured Analyzer-to-Refiner workflow. By separating evidence-grounded diagnosis from targeted refinement, the modular design supports a more interpretable and controllable process and leads to more consistent gains across datasets and model backbones.
5.3 Qualitative analysis
We additionally conducted a human evaluation to examine two complementary aspects of MA-SAPO. [^8] The full evaluation protocol, per-dimension scores, and the rating distribution are provided in Appendix F. A total of 17 annotators, comprising data scientists, AI engineers, and graduate students in data science or related fields, participated in the study. For reasoning quality (H1), MA-SAPO achieved a mean score of 3.93 on a 5-point scale, outperforming the single-agent baseline by 0.24 points, with notable gains in usefulness and accuracy. For directional consistency (H2), annotators assessed whether optimized prompts retained the original intent, yielding a mean rating of 3.36 out of 4.0, indicating that prompt optimization largely preserves semantic alignment. Taken together, the human evaluation supports the quantitative findings and demonstrates that MA-SAPO improves output quality without distorting the original task intent.
5.4 Cost and efficiency analysis
Figure 3: Efficiency–performance trade-off of MAD, MARS, and MA-SAPO. Panel (a) plots average input tokens per instance against average score, and panel (b) plots execution time per instance against average score, under GPT-4o and Llama3-8B. Averaged over HelpSteer1 and HelpSteer2, MA-SAPO delivers the highest scores with the lowest token cost and substantially lower latency than MARS.
To complement the effectiveness results, we also compare the computational cost of different multi-agent frameworks. Figure 3(a) reports the average input tokens per instance against average score. MAD incurs moderate overhead but yields relatively weak overall performance, while MARS improves performance at a substantially higher token cost, consuming over 17K input tokens per instance. In contrast, MA-SAPO reduces token consumption to approximately 5K input tokens per instance by shifting reasoning asset construction offline and using a single Analyzer-to-Refiner loop at test time, while achieving the highest average scores under both GPT-4o and Llama3-8B. Figure 3(b) further illustrates the trade-off between execution time per instance and average score. MAD achieves the shortest execution time at approximately 5 seconds per instance, yet delivers the lowest average performance among the three methods. MARS incurs the highest latency, requiring roughly 85 to 90 seconds per instance. By contrast, MA-SAPO maintains an execution time of approximately 40 seconds per instance, which is less than half that of MARS, while delivering substantially higher scores. Taken together, these results show that MA-SAPO achieves the most favorable efficiency-performance balance across both token cost and latency. [^9] Detailed model inference hardware specifications are provided in Appendix D.
6 Conclusion
We propose MA-SAPO, a novel multi-agent framework for interpretable and controllable prompt optimization. Compared to prior methods that either perform one-shot rewrites, rely on unguided retrieval, or employ heavy multi-agent debates, MA-SAPO explicitly transforms evaluation outcomes into reusable reasoning assets and grounds refinements in evidence-backed Analyzer-to-Refiner interactions. [^10] Practical implications are discussed in Appendix G. Our contributions extend beyond framework design to include comprehensive experiments on two benchmarks and multiple backbones, ablation studies on retrieval depth and agent configuration, and human evaluations of reasoning quality and semantic consistency. Extensive results demonstrate that MA-SAPO consistently outperforms existing approaches, achieving higher optimization quality while remaining efficient in cost and latency.
Limitations
Although MA-SAPO achieves superior performance compared to a range of single-agent reasoning baselines and interaction-based agent frameworks, these results do not imply that MA-SAPO uniformly dominates all baselines across tasks and conditions. MA-SAPO leverages retrieved exemplars as evidence for diagnosis and prompt rewriting. However, retrieval may suffer from limited recall under paraphrase or lexical mismatch, causing the system to rely on less relevant assets and creating a bottleneck at the retrieval stage. Moreover, the single-path, sequential multi-agent pipeline is prone to cascading errors and provides limited exploration, since it does not explicitly branch over, compare, and select among alternative diagnostic/edit trajectories. Future work could improve robustness by adding verifier-based artifact validation, uncertainty-aware gated updates, and constrained rewriting, as well as by introducing multi-path candidate generation with iterative verification.
Appendix
Appendix A Datasets detail
@@
We conduct our experiments on the HelpSteer family of datasets (Wang et al., 2024c; b), which provide human-annotated prompt-response pairs for evaluating prompt optimization in large language models. Both datasets share the same annotation schema across five quality dimensions, but differ in size, design, and annotation depth, as summarized in Table 4. We build our retrieval corpus from the HelpSteer2 training set and conduct evaluation on the validation splits of both datasets.
@@
Complexity Intellectual depth required to write the response.
@@
Table 4: Overview of HelpSteer1/2 annotation dimensions. Each scored on a 0–4 scale.
Figure 4: Score distributions across five annotation dimensions for HelpSteer1 (left) and HelpSteer2 (right). Each bar shows the proportion of samples assigned each score from 0 to 4. Coherence is heavily concentrated at higher scores in both datasets, while complexity and verbosity exhibit broader distributions across the full range.
A.1 HelpSteer1
HelpSteer1 contains 35.3k training and 1.79k validation samples, each comprising a prompt, a response, and five human-annotated attributes scored on a 0–4 scale. Prompts cover a broad range of task types including rewriting, summarization, classification, extraction, and question answering, with up to four responses generated per prompt. Each response was independently reviewed by multiple annotators, providing reliable quality estimates across diverse prompt-response styles.
A.2 HelpSteer2
HelpSteer2 is a newer dataset containing 20.3k training and 1.04k validation samples, with higher-quality annotations, multi-turn prompts (about 29% of cases), and additional preference annotations between responses (Wang et al., 2024b). Compared to HelpSteer1, it employs a substantially larger annotator pool and a more diverse set of response sources, with each sample receiving a minimum of three independent annotations. We adopt it as our retrieval corpus since (i) its richer annotations and multi-turn coverage yield more representative reasoning assets, (ii) preference signals provide finer-grained evidence for our Analyzer and Refiner agents, and (iii) it is the current benchmark standard for training and evaluating reward models, ensuring alignment with best practices in the field. As shown in Figure 4, HelpSteer2 exhibits more concentrated distributions toward higher scores, particularly in coherence, reflecting the higher quality of responses.
Appendix B Baseline details